Minutes of the RSC Meeting on Friday, April 18th

From: Flemming Videbaek (videbaek@sgs1.hirg.bnl.gov)
Date: Tue Apr 22 2003 - 17:12:21 EDT

  • Next message: Flemming Videbaek: "pp run"
    In addition to the RHIC coordination meetings the RSC* has a weekly meeting (Friday afternoon's late)
    to discuss progress and issues in relation to the pp spin runnning.
    
    Brendan Fox, who is in the Riken Spin group, is working with us to help out on the transverse asymmtry measurements
    in particular how to measure the luminosities of the different spin-states in the beam using scalers, and how to use this
    information to extract and evaluate errors on the (to be measured) assymmetry.
    
    He has written up the following minutes from the Friday meeting.
    
    regards
        Flemming
    
    
    RSC* Rhic Spin Collaboration
    An umbrella group for the spin interests of all experiments.
    ------------------------------------------------------
    Flemming Videbaek
    Physics Department
    Brookhaven National Laboratory
    
    tlf: 631-344-4106
    fax 631-344-1334
    e-mail: videbaek@bnl.gov
    
    ----- Original Message -----
    From: "Brendan Fox" <fox@drax.Colorado.EDU>
    To: "Spin Commissioning List" <spin-commission-l@lists.bnl.gov>
    Cc: "Star Spin List" <starspin-hn@connery.star.bnl.gov>; "Phenix Spin List" <phenix-spin-l@bnl.gov>; "Spin Discussion
    Email List" <spin-discussion-l@bnl.gov>; "Wit Busza" <busza@mit.edu>; "Mark Baker" <baker@rcf.rhic.bnl.gov>; "Gunther
    Roland" <gunther.roland@cern.ch>; "Dave Hofman" <hofman@uic.edu>; "Flemming Videbaek" <videbaek@rcf.rhic.bnl.gov>; "Dana
    Beavis" <beavis@bnl.gov>
    Sent: Tuesday, April 22, 2003 1:56 PM
    Subject: Minutes of the RSC Meeting on Friday, April 18th
    
    
    |
    | Hi y'all,
    |
    | Here is the summary of last Friday's RSC meeting ...
    |
    |                                                                 ... brendan
    |
    | ============================================================================
    |
    |
    |  (1) Status of RHIC Commissioning [Haixin]
    |
    |      what has been happened over the last week?
    |
    |      - have ramped 55x55 with beta*=1m & 0.5*10^11 particles per bunch.
    |      - with 110x110 with 10^{11} particles per bunch, saw significant pressure
    |        rises; would take 2+ days to scrup the vacuum chambers with the beam
    |        to solve the problem.
    |      - measured polarization in both blue and yellow at flat-top
    |        * blue tends to be in 25% to 35% range
    |        * yellow tends to be around 20%; the difference may be coming from
    |          the tune crossing which happens during the beta* squeeze since the
    |          tune gets close to the 3/14 resonance.
    |      - PLL works and is presently being used for tune-development.
    |      - one of the inner helices in one of the yellow snakes died.
    |      - aperture scan on yellow, learned that vertical bump in yellow to
    |        need for accomodating the plan for running the damaged snake is ~1.5cm.
    |
    |      what is planned?
    |
    |      - focus on 55x55 bunch mode so that we can get to physics running soon.
    |      - implement new procedures to prevent snake quenches.
    |      - develop new ramp with only the outer helices of the damaged yellow
    |        snake; it is hoped that this can be accomplished over the weekend
    |      - study the loss of polarization in yellow during the ramp.
    |      - continue with the spin rotator commissioning.
    |
    |      as far as timeline,
    |
    |      April 18-20 : setup new ramp; get collisions by Monday
    |            21-22 : continue to commissioning the spin rotators
    |              23  : access
    |            24-25 : continue to commissioning the spin rotators
    |            26-27 : maximize intensity for 55x55 mode
    |
    |  (2) Status of the Yellow Snake [Thomas+Alfredo]
    |
    |      what happened?
    |
    |      - the inner helix of one of the yellow snakes was damaged during a
    |        quench and cannot be repaired or used.
    |
    |      what is the solution?
    |
    |      - use only the outer helices of the damaged snake.  To do so, it
    |        is necessary to:
    |
    |        * run the outer coils at 300A instead of 100A.  This is not a
    |          problem since all snake coils were tested for running at 350A.
    |        * the vertical excursion in the snake is now about 1 cm larger
    |          at injection; so the bumps need to be adjusted at the snake so
    |          that the beam goes through it.
    |        * live with the fact that the stable direction of the polarization
    |          is now about 11 degrees off the vertical.  This results in a
    |          loss of polarization of 2%.  Spin tracking, as shown by Alfredo,
    |          showed that a large emittance particle seemed to have ~96%
    |          polarization and thus confirmed the calculation.
    |        * radial component of the polarization will depend on the position
    |          in the ring.
    |        * could increase the emittance dependence of the polarization a bit.
    |
    |      what needs to be done to implement this solution?
    |
    |      - need slightly better tune control
    |      - retune the ramp with the new snake configuration.
    |      - measure the value and direction of the polarization using the
    |        RHIC CNI polarimeter.
    |      - try to compensate for the 11 degree tilt using the spin rotators.
    |
    |  (3) Status of the Rotator Commissioning [Waldo]
    |
    |      what has been accomplished?
    |
    |      - on Tuesday, after 3 or 4 tries with only beam in blue (due to the
    |        damaged snake in yellow), got 6x6 bunches to flat-top with the
    |        blue rotators at PHENIX at a ramp rate of 1/4 A/s.  Saw that the
    |        beam lifetime would drop when ramping the rotator, returned to
    |        normal after ramping.
    |
    |      what needs to be done?
    |
    |      - finalize ramp rate; Waldo thinks that 1/8 A/s would be better but
    |        then it would take 27 minutes to bring up the rotators. Some
    |        more learning here as to what is doable.
    |      - turn on yellow rotators.
    |      - check beam polarization w/rotators on using RHIC CNI polarimeters.
    |      - check effect of spin rotators with PHENIX local polarimeter, hoping
    |        to do this on Monday.
    |      - think about whether the rotators can compensate for the 11 degree
    |        rotation of the stable spin axes.
    |
    |  (4) Status of the AGS CNI Polarimeter [Jeff]
    |
    |      what has been accomplished?
    |
    |      - one of the great success of this week is that the AGS CNI polarimeter
    |        has successfully measured the beam polarization along the ramp.
    |        In these data, one sees:
    |        * the spin flips at each imperfection resonances
    |        * the raw asymmetry value falls from 0.008 to 0.004
    |        * there appears to be some steps at 36-Qy, 24+Qy, 48-Qy, and 36+Qy
    |          resonances.  More statistics would be helpful here.
    |      - they have data at G-gamma=12.5 (beam momentum of ~6 GeV/c) for both
    |        the CNI and E880 polarimeters.  The understanding of these data is
    |        still in progress.
    |
    |      what is planned?
    |
    |      - continue to measure the polarization along the ramp to improve the
    |        statistics and thus help discern steps at intrinsic resonances
    |      - cross calibrate the AGS CNI with the E880 polarimeter
    |
    |  (5) Status of the AGS Commissioning [Leif]
    |
    |      what was accomplished?
    |
    |      - decreased the emittances by:
    |        * (H)   improving the optical matching between the booster & AGS
    |        * (V,L) scraping the beam in the booster now that the source is
    |                delivering more beam
    |      - with the lower emittance, the polarization is now in the 40-45%
    |        range when the source polarization was 65% or so.
    |      - increased the source polarization from 65% to 75%.
    |
    |      what is planned?
    |
    |      - continue to look at the polarization along the ramp using the CNI
    |
    |  (6) Status of the RHIC CNI Polarimeters [Osamu]
    |
    |      what has been accomplished?
    |
    |      - target are now withdrawn before DAQ readout
    |      - transition between "AT" event-by-event & "scaler" mode operation now
    |        implemented in control software
    |      - banana cut has been fine-tuned
    |      - QA/offline analysis code now ready for "AT" event-by-event mode
    |      - developed quality plots, looking at ratio of # of carbon events to
    |        total events and # of carbon events to background events
    |      - development of "on the ramp" measurement, have learned:
    |        * see item 15 below for comments on rate/quality of measurement
    |        * The time shift is 7 to 8 ns along the ramp, small compared to the
    |          timing window, so standard window can be used.
    |        * Target position be the same during the ramp since the beam movement
    |          is ~1 mm over the course of the ramp.
    |        * Energy information can be determined by sync'ng with the RHIC
    |          clock.  Expect to put in delimiters at 20 Hz.
    |        * Can adjust the carbon banana cut as needed in offline analysis since
    |          the data will be collected in "AT" event-by-event mode.
    |
    |      what is planned?
    |
    |      - at injection, take scaler mode data in first filled beam, then
    |        event-by-event mode data in second filled beam.  This difference
    |        arises from the ~5 minute readout time per "AT" mode measurement and
    |        the desire to start ramping the beams as soon as possible.  In this
    |        way, the ~5 minute readout time for the second measurement overlaps
    |        with ramping.
    |      - take "AT" mode data in both rings at flattop
    |      - continue to develop "on the ramp" measurement; make a measurement on
    |        the ramp.
    |
    |  (7) Status of PHENIX [Abhay]
    |
    |      what has been accomplished?
    |
    |      - shutdown work was successful
    |      - DAQ work continues
    |      - CDEV data has been incorporated into the PHENIX datastream
    |
    |      what is planned?
    |
    |      - commission the spin rotators with PHENIX local polarimeter
    |      - study the relative luminosity to determine the size of the systematic
    |        error
    |
    |  (8) Status of STAR [Les]
    |
    |      what has been accomplished?
    |
    |      - BBC commissioning done/bad scaler channels were fixed during the access
    |      - STAR ready for data
    |      - east FPD, all ADCs installed, partially commissioned
    |
    |      what is planned?
    |
    |      - complete commissioning of FPDs
    |      - final commissioning of the BBC by looking at polarization effects
    |      - commission the jet trigger
    |
    |      and then some questions ...
    |
    |      Will tuning PHENIX rotators effect STAR collisions/background? [Les]
    |
    |      Waldo doesn't think so, but will not really know until he sees some
    |      data during the upcoming week.
    |
    |  (9) Status of BRAHMS [Brendan]
    |
    |      Waiting for one night of good luminosity beam and then will start to
    |      take survey data.  Having tested the spin-sorted scalers, they have
    |      been moved to the IR and now we are working on setting them up.
    |
    | (10) Status of pp2pp [Wlodek]
    |
    |      What has been accomplished?
    |
    |      - during the 2-day access, they installed all of their Roman pots and
    |        the necessary readout electronics
    |      - the trigger scintillators have been checked out
    |
    |      What needs to be done?
    |
    |      - finish work on the DAQ
    |      - test the silicon strip detectors in situ using random triggers (mostly,
    |        looking at the noise level here).
    |      - hopefully ready for beam at the end of next week.
    |
    | (11) Discussion of the Spin Pattern [Gerry]
    |
    |      The choices:
    |
    |      (a) use one pattern   -- namely, the one proposed by Waldo
    |      (b) use two patterns  -- Waldo's + variant of Waldo's proposal
    |      (c) use four patterns -- Waldo's + three variants of Waldo's proposal
    |      (d) use random pattern
    |
    |      Decision: (c), the four patterns, an example set being:
    |
    |      (1) B +-+-  [Waldo's original proposal]
    |          Y ++--
    |
    |      (2) B ++--  [Waldo's original proposal + switch the beams]
    |          Y +-+-
    |
    |      (3) B +-+-  [Waldo's original proposal + flip one of the patterns]
    |          Y --++
    |
    |      (4) B --++  [Waldo's original proposal + switch the beams
    |          Y +-+-                             & flip one of the patterns]
    |
    |      Gerry wishes to think a bit more about the exact pattern choice to
    |      see if there isan option which better balances parallel and
    |      anti-parallel.
    |
    | (12) Discussion of the Spin Flipper [Gerry]
    |
    |      Since time is precious, how should we proceed with the spin flipper?
    |
    |      Proposal:
    |
    |      In addition to reducing the relative luminosity error significantly,
    |      the spin flipper also allows us to learn a bit about the spin tune.
    |      So, it was felt that the reasonable approach would be:
    |
    |      (a) begin commissioning work at the end of stores.  In this way, we
    |          learn how it works.
    |      (b) given the new conditions with beta*=1m at both PHENIX and STAR, it
    |          is not wise to extrapolate last year's performance to this year.
    |          So, the experiments should report on the systematic error in the
    |          relative luminosity measurement as soon as possible.  If this
    |          error is already small enough, we can live without the spin
    |          flipper during the physics run.  If not, we need it.
    |
    |      In this way, we follow the suggestion of Les to focus on getting the
    |      physics run underway now and, as necessary and time permitting, address
    |      other issues.  So, we will revisit this discussion next Friday when we
    |      have more information.
    |
    | (13) Discussion of the Downramp [Gerry]
    |
    |      Due to limited tries, there is some likelihood that the downramp will
    |      not be accomplished this year.  At this time, it was felt that it was
    |      premature to have a lengthy conversation on this matter.  For now, we
    |      will focus on getting the physics run underway.  But, we should be
    |      prepared to consider the downramp later if time permits.
    |
    | (14) Discussion of Longitudinal Bunch Squeezing [Brendan]
    |
    |      - with the narrower emittance bunches from the AGS, PHENIX say that
    |        the ratio of collisions w/in the acceptance to all collisions
    |        was as high as 60% (compared to 20 to 30% nominally) with 55x55
    |        and intensities of 0.5*10^11/bunch.
    |      - Mike Brennen believes that the rF squeezing should be possible; just
    |        needs some time try it.  Haixin says that, at some point during the
    |        upcoming week, it will be tried.
    |
    | (15) Discussion on Measuring the Polarization Along the RHIC Ramp [Hal/Gerry]
    |
    |      Hal asked about the statistics of the proposed measurement of the
    |      polarization along the ramp.  During the meeting, the answer was not
    |      quantatative, so Gerry sat down and worked out the numbers.
    |
    |      "The RHIC ramp takes ~2 minutes (compared to ~0.5s in the AGS) and the
    |       polarization direction is unchanged in RHIC (compared to the flipping
    |       every 523 MeV in the AGS).  So, the data can be binned into just a few
    |       bins to test certain resonances and the beta* squeeze.  In RHIC, there
    |       are 55 bunches (compared to 3 in the AGS during the ramp studies).  The
    |       detectors in RHIC are ~3/5 closer the beam than those in the AGS and there
    |       are two 45 degree detectors in RHIC, so the effective acceptance for a
    |       vertical polarization is about a factor of 4 larger in RHIC than in the
    |       AGS.  The RHIC target is approximately 1/100 thinner than the AGS target.
    |       We expect to collected 50M events in 1 ramp in RHIC.  If five bins are
    |       used, the error on the raw asymmetry will be \sqrt{10^7} or 3*10^{-4}.
    |       For a polarization of 0.4 and A_N=0.01, this precision gives a measurement
    |       of \delta{P}/P=8% for each bin (\delta{P}=0.032).  So, with only one
    |       ramp measurement, we would be able to locate a polarization loss of 0.07
    |       or larger."
    |
    


    This archive was generated by hypermail 2.1.5 : Tue Apr 22 2003 - 17:09:02 EDT